OpenStudy (anonymous):

-3x^4+27x^2+1200=0 find all zeros of the equation! PLEASE HELP! I will medal

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@ganeshie8

OpenStudy (muzzack):

3

OpenStudy (muzzack):

real solutions are 5 and -5

OpenStudy (anonymous):

i need work to show

OpenStudy (anonymous):

please show work. help me thru it

OpenStudy (anonymous):

@helpme1.2

OpenStudy (anonymous):

If you have a graphing calculator...you could draw a picture, and then to calculate the intercepts, press 2nd trace zero

OpenStudy (anonymous):

but i have to show written work

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@ganeshie8 please help

OpenStudy (factor):

wht is ^ this symbol

OpenStudy (anonymous):

@Muzzack please help

OpenStudy (anonymous):

means square of some number

OpenStudy (muzzack):

okay well first try to make the x's add or subtract by doing this leave the numbers multiplying them dividing to make the x's alone, and we know what one thing did in one side must do to the other

OpenStudy (muzzack):

still get it

OpenStudy (anonymous):

show work, i dont understand D:

OpenStudy (anonymous):

HELLO????? @Muzzack

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

what i got for this problem is x=25,-16 is that even right????

OpenStudy (anonymous):

i just factored.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

\[-3x ^{4}+27x ^{2}+1200=0 Divide all of this by -3 \]

OpenStudy (anonymous):

\[x ^{4}-9x ^{2}-400\] FACTOR THIS

OpenStudy (anonymous):

(x-25)(x+16) =0 x=25,-16

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

uhh idk how to get that

OpenStudy (anonymous):

can you put it in one message? im so confused

OpenStudy (anonymous):

not my factored one no